# Deep SPI: Safe Policy Improvement via World Models

### v1

@@ Abstract
Safe policy improvement (SPI) offers theoretical control over policy updates, yet existing guarantees largely concern offline, tabular reinforcement learning (RL). We study SPI in general online settings, when combined with world model and representation learning. We develop a theoretical framework showing that restricting policy updates to a well-defined neighborhood of the current policy ensures monotonic improvement and convergence. This analysis links transition and reward prediction losses to representation quality, yielding online, “deep” analogues of classical SPI theorems from the offline RL literature. Building on these results, we introduce DeepSPI, a principled on-policy algorithm that couples local transition and reward losses with regularised policy updates. On the ALE-57 benchmark, DeepSPI matches or exceeds strong baselines, including PPO and DeepMDPs, while retaining theoretical guarantees.
@@ 1 Introduction
SPI methods provide principled guarantees on policy updates from fixed datasets (offline RL) (Thomas et al., 2015; Ghavamzadeh et al., 2016a; Laroche et al., 2019; Simão et al., 2020; Castellini et al., 2023). These methods assume tabular state spaces and offline data, where error bounds must hold globally across all state–action pairs, often via robust MDP formulations (Iyengar, 2005; Nilim and Ghaoui, 2005). Our setting is fundamentally different: we study online RL with high-dimensional inputs, where such global constraints are intractable. We take inspiration from the SPI literature but introduce local, on-policy losses that make safe improvement feasible in practice.
Representation learning and model-based RL. Auxiliary transition and reward prediction losses are central to many model-based methods, from DeepMDP to Dreamer and related world-model approaches (Gelada et al., 2019; Hafner et al., 2021). In particular, the losses we consider for learning transitions and rewards generalize a wide range of objectives used across the model-based RL literature (François-Lavet et al., 2019; van der Pol et al., 2020; Kidambi et al., 2020; Delgrange et al., 2022; Dong et al., 2023; Alegre et al., 2023). Other works design representations that cluster states into groups within which the agent is guaranteed to behave similarly (Zhang et al., 2021; Castro et al., 2021; Agarwal et al., 2021; Avalos et al., 2024), typically under restrictive conditions (e.g., deterministic dynamics or Gaussian-kernel assumptions). By contrast, we directly link representation quality and model accuracy to our safe policy improvement analysis, yielding tractable guarantees in the online setting.
@@ 2 Background
At each time step $t$, the agent selects an action according to a (stationary) policy $\pi:\mathcal{S}\to\Delta\mathopen{}\mathclose{{\left(\mathcal{A}}}\right)$ mapping states to distributions over actions. Running an MDP under $\pi$ induces a unique probability measure $\mathbb{P}_{\pi}$ over trajectories (Revuz, 1984), with associated expectation operator $\mathbb{E}_{\pi}$; we write $\mathbb{E}_{\pi}[\cdot\mid s_{0}=s]$ when the initial state is fixed to $s\in\mathcal{S}$. A policy has full support if $\text{supp}\mathopen{}\mathclose{{\left(\pi(\cdot\mid s)}}\right)=\mathcal{A}$ for all $s\in\mathcal{S}$, and we denote the set of all policies by $\Pi$. The stationary measure of $\pi$ is the distribution over states visited under $\pi$, defined by $\xi_{\pi}(\cdot)=\mathbb{E}_{s\sim\xi_{\pi}}\mathbb{E}_{a\sim\pi(\cdot\mid s)}[P(\cdot\mid s,a)]$. This measure is often assumed to exist in continual RL (Sutton and Barto, 2018), is unique in episodic RL (Huang, 2020), and corresponds to the occupancy measure in discounted RL (Metelli et al., 2023). [^1] Details on the formalization of episodic processes and value functions can be found in Appendix A.
@@
Representation learning in RL. In realistic environments, the state–action space is too large for tabular policies or value functions. Instead, deep RL employs an encoder $\phi\colon\mathcal{S}\to\,\overline{\!{\mathcal{\mathcal{S}}}}$ that maps states to a tractable latent space $\overline{\!{\mathcal{\mathcal{S}}}}$, from which value functions can be approximated. Learning such encoders is referred to as representation learning. To improve representations, agents are often trained with additional objectives, commonly auxiliary tasks requiring predictive signals. Policy-based methods then optimize a latent policy $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\colon\,\overline{\!{\mathcal{\mathcal{S}}}}\to\Delta\mathopen{}\mathclose{{\left(\mathcal{A}}}\right)$ jointly with $\phi$, executed in the environment as $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu(\cdot\mid\phi(s))$. By convention, we write $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu(\cdot\mid s)$ for $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi(s)$ when $\phi$ is clear, and denote the set of all latent policies by $\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$. For any $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\in\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$, the composed policy $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi$ belongs to $\Pi$.
Model-based RL augments policy learning with a world model $\overline{\!{\mathcal{M}}}=\langle{\,\overline{\!{\mathcal{\mathcal{S}}}},\mathcal{A},\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu,\,\overline{\!{{R}}},\bar{s}_{I},\gamma\rangle}$, which can improve (i) sample efficiency by generating trajectories, (ii) value estimation through planning, and (iii) representation learning by grouping states with similar behavior. When $\overline{\!{\mathcal{\mathcal{S}}}}=\mathcal{S}$, the model must replicate environment dynamics, which is often intractable. Instead, we focus on $\overline{\!{\mathcal{\mathcal{S}}}}$ defined by the learned representation $\phi$, so that $\overline{\!{\mathcal{M}}}$ becomes an abstraction of $\mathcal{M}$. Learning transition and reward functions then additionally serves as an auxiliary signal for the representation, encouraging states with similar behavior to map close in $\overline{\!{\mathcal{\mathcal{S}}}}$. Since $\overline{\!{\mathcal{\mathcal{S}}}}$ is the latent space, $\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$ corresponds to the policies of $\overline{\!{\mathcal{M}}}$. We further assume $\overline{\!{\mathcal{\mathcal{S}}}}$ is equipped with a metric $\overline{\!{d}}\colon\,\overline{\!{\mathcal{\mathcal{S}}}}\times\,\overline{\!{\mathcal{\mathcal{S}}}}\to[0,\infty)$ to measure distances.
@@ 3 No way Home: when world models and policies go out of trajectories
World models are usually learned toward minimizing a reward loss $L_{{R}}$ and/or transition loss $L_{P}$ from experiences $\eta$ collected along the agent’s trajectories. Those experiences are either gathered in the form of a batch or a replay buffer $\mathcal{B}$. In general, the loss functions take the following form: $L_{{R}}=\mathop{\mathbb{E}}_{\eta\sim\mathcal{B}}f_{{R}}\mathopen{}\mathclose{{\left(\phi,\,\overline{\!{{R}}};\,\eta}}\right)$ and $L_{P}=\mathop{\mathbb{E}}_{\eta\sim\mathcal{B}}f_{P}\mathopen{}\mathclose{{\left(\phi,\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu;\,\eta}}\right)$, where $f_{R}$ (resp. $f_{P}$) assign a “cost” relative to the error between R{R} and $\overline{\!{{R}}}$ (resp. $P$ and $\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu$) according to the experiences $\eta$ and their representation. Henceforth, we refer to the policy $\pi_{\text{b}}$ used to insert experiences in $\mathcal{B}$ as the baseline policy.
@@ 3.1 Out-of-trajectory world model
To illustrate this problem, consider the world model of Figure 1. Assume the model is trained by collecting trajectories produced by $\pi_{\text{b}}$ in $\mathcal{M}$ where $\pi_{\text{b}}\mathopen{}\mathclose{{\left(a_{2}\mid s}}\right)\leq\epsilon$ for all $s\in\mathcal{S}_{1}$, with $\epsilon>0$. For a sufficiently small $\epsilon$, the region $\mathcal{S}_{3}$ in the original environment would remain largely unexplored while having almost no impact on the losses $L_{{R}},L_{P}$. Therefore, the representation of states in $\mathcal{S}_{3}$ ($\bar{s}_{3}$ and $\bar{s}^{\prime}_{3}$) may turn completely inaccurate. Here, the model incorrectly assigns a reward of $20$ to $\bar{s}^{\prime}_{3}$, whereas the true reward is strictly negative. Consequently, the optimal policy in $\overline{\!{\mathcal{M}}}$ deterministically selects $a_{2}$ in $\bar{s}_{1}$. When executed in the original environment, this policy drives the agent to $\mathcal{S}_{3}$ thereby degrading the baseline policy $\pi_{\text{b}}$.
@@ 3.2 Confounding policy update
Updating both the representation and the policy solely from experience collected under a baseline policy can degrade performance rather than improve it. In the same spirit as policy confounding (Suau et al., 2024), we call this phenomenon confounding policy update. The MDP in Figure 2 illustrates the issue.
@@
The agent maps the states $s_{2}$ and $s_{3}$ to the same latent state $\bar{s}$, i.e. $\phi(s)=\bar{s}$ iff $s\in\{s_{2},s_{3}\}$. States $s_{1}$ and $s_{4}$ each have their own latent state. We consider the baseline policy $\pi_{\text{b}}:=\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}\circ\phi$, where $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}$ is a stochastic policy with a small exploration rate $\zeta$:
@@
for $0<\zeta\ll\epsilon$. Ideally, a good representation would group states from which the agent behaves similarly. Because trajectories that reach $s_{3}$ and pick $a_{2}$ are unlikely, the two states appear identical under $\pi_{\text{b}}$: $|V^{\pi_{\text{b}}}(s_{2})-V^{\pi_{\text{b}}}(s_{3})|\!\approx\!0$. Therefore, this justifies using $\phi$ as representation for $\pi_{\text{b}}$, since the values of $s_{2}$ and $s_{3}$ are nearly identical — the agent exhibits close behaviors under $\pi_{\text{b}}$ from those states.
@@
Our goal is to establish sufficient conditions to guarantee safe policy improvement during the RL process, either based on world models, state representations, or both, thus alleviating OOT world model and confounding policy update issues. Notice that, in the examples, both problems occur when performing aggressive updates from $\pi_{\text{b}}$ to a new policy $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ (the mode of the distributions drastically shifts). Intuitively, smooth updates indeed ensure to alleviate those issues: constraining the policy search to policies “close” to $\pi_{\text{b}}$ (i) prevents hallucinations in parts of the world model that have been underexplored; (ii) reduces the risk of significantly degrading the return when updating the policy. While the benefits of regularizing policy improvements has been already both theoretically and practically justified (e.g., Geist et al. 2019), their implications when mixing model-based and representation learning in RL has been underexplored.

### v2

@@ Abstract
Safe policy improvement (SPI) offers theoretical control over policy updates, yet existing guarantees largely concern offline, tabular reinforcement learning (RL). We study SPI in general online settings, when combined with world model and representation learning. We develop a theoretical framework showing that restricting policy updates to a well-defined neighborhood of the current policy ensures monotonic improvement and convergence. This analysis links transition and reward prediction losses to representation quality, yielding online, “deep” analogues of classical SPI theorems from the offline RL literature. Building on these results, we introduce DeepSPI, a principled on-policy algorithm that couples local transition and reward losses with regularised policy updates. On the ALE-57 benchmark, DeepSPI matches or exceeds strong behaviorals, including PPO and DeepMDPs, while retaining theoretical guarantees.
@@ 1.1 Related Work
SPI methods provide principled guarantees on policy updates from fixed datasets (offline RL) (Thomas et al., 2015; Ghavamzadeh et al., 2016a; Laroche et al., 2019; Simão et al., 2020; Castellini et al., 2023). These methods assume tabular state spaces and offline data, where error bounds must hold globally across all state–action pairs, often via robust MDP formulations (Iyengar, 2005; Nilim and Ghaoui, 2005). Our setting is fundamentally different: we study online RL with high-dimensional inputs, where such global constraints are intractable. We take inspiration from the SPI literature but introduce local, on-policy losses that make safe improvement feasible in practice. In spirit, other model-based methods share the goal of providing SPI-like guarantees in more general settings, but are again purely offline, omit any form of representation learning, and rely on assumptions that differ substantially from ours (Yu et al., 2020; 2021; Kidambi et al., 2020).
Representation learning and model-based RL. Auxiliary transition and reward prediction losses are central to many model-based methods, from DeepMDP to Dreamer and related world-model approaches (Gelada et al., 2019; Hafner et al., 2021). In particular, the losses we consider for learning transitions and rewards generalize a wide range of objectives used across the model-based RL literature (François-Lavet et al., 2019; van der Pol et al., 2020; Kidambi et al., 2020; Delgrange et al., 2022; Dong et al., 2023; Alegre et al., 2023). Conceptually, our representation guarantees are closely connected to classical notions of state abstraction in MDPs (Li et al., 2006) and to bisimulation (Larsen and Skou, 1991; Desharnais et al., 1998). Building on bisimulation, prior work develops representations that group states into areas in which the agent is guaranteed to behave similarly under the current policy (Castro, 2020; Zhang et al., 2021; Castro et al., 2021; Agarwal et al., 2021a; Avalos et al., 2024). By contrast, we directly link representation quality and model accuracy to our safe policy improvement analysis, yielding tractable guarantees in the online setting.
@@ 2 Background
At each time step $t$, the agent selects an action according to a (stationary) policy $\pi:\mathcal{S}\to\Delta\mathopen{}\mathclose{{\left(\mathcal{A}}}\right)$ mapping states to distributions over actions. Running an MDP under $\pi$ induces a unique probability measure $\mathbb{P}_{\pi}$ over trajectories (Revuz, 1984), with associated expectation operator $\mathbb{E}_{\pi}$; we write $\mathbb{E}_{\pi}[\cdot\mid s_{0}=s]$ when the initial state is fixed to $s\in\mathcal{S}$. A policy has full support if $\text{supp}\mathopen{}\mathclose{{\left(\pi(\cdot\mid s)}}\right)=\mathcal{A}$ for all $s\in\mathcal{S}$, and we denote the set of all policies by $\Pi$. A stationary measure of $\pi$ is a distribution over states visited under $\pi$, and is defined as a solution of $\xi_{\pi}(\cdot)=\mathbb{E}_{s\sim\xi_{\pi}}\mathbb{E}_{a\sim\pi(\cdot\mid s)}[P(\cdot\mid s,a)]$. Such a measure is often assumed to exist in continual RL (Sutton and Barto, 2018), is unique in episodic RL (Huang, 2020), and defines the occupancy measure in discounted RL (Metelli et al., 2023). [^1] Details on the formalization of episodic processes and value functions can be found in Appendix A.
@@
Representation learning in RL. In realistic environments, the state–action space is too large for tabular policies or value functions. Instead, deep RL employs an encoder $\phi\colon\mathcal{S}\to\,\overline{\!{\mathcal{\mathcal{S}}}}$ that maps states to a tractable latent space $\overline{\!{\mathcal{\mathcal{S}}}}$, from which value functions can be approximated. Learning such encoders is referred to as representation learning (Echchahed and Castro, 2025). To improve representations, agents are often trained with additional objectives, commonly auxiliary tasks requiring predictive signals. Policy-based methods then optimize a latent policy $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\colon\,\overline{\!{\mathcal{\mathcal{S}}}}\to\Delta\mathopen{}\mathclose{{\left(\mathcal{A}}}\right)$ jointly with $\phi$, executed in the environment as $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu(\cdot\mid\phi(s))$. By convention, we write $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu(\cdot\mid s)$ for $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi(s)$ when $\phi$ is clear, and denote the set of all latent policies by $\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$. For any $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\in\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$, the composed policy $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi$ belongs to $\Pi$.
Model-based RL augments policy learning with a world model $\overline{\!{\mathcal{M}}}=\langle{\,\overline{\!{\mathcal{\mathcal{S}}}},\mathcal{A},\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu,\,\overline{\!{{R}}},\bar{s}_{I},\gamma\rangle}$, which can improve (i) sample efficiency by generating trajectories (e.g., Hafner et al. 2021), (ii) value estimation through planning (e.g., Buckman et al. 2018), and (iii) representation learning by grouping states with similar behavior (e.g., Gelada et al. 2019; Zhang et al. 2021). When $\overline{\!{\mathcal{\mathcal{S}}}}=\mathcal{S}$, the model must replicate environment dynamics, which is often intractable. Instead, we focus on $\overline{\!{\mathcal{\mathcal{S}}}}$ defined by the learned representation $\phi$, so that $\overline{\!{\mathcal{M}}}$ becomes an abstraction of $\mathcal{M}$. Learning transition and reward functions then additionally serves as an auxiliary signal for the representation, encouraging states with similar behavior to map close in $\overline{\!{\mathcal{\mathcal{S}}}}$. Since $\overline{\!{\mathcal{\mathcal{S}}}}$ is the latent space, $\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$ corresponds to the policies of $\overline{\!{\mathcal{M}}}$. We further assume $\overline{\!{\mathcal{\mathcal{S}}}}$ is equipped with a metric $\overline{\!{d}}\colon\,\overline{\!{\mathcal{\mathcal{S}}}}\times\,\overline{\!{\mathcal{\mathcal{S}}}}\to[0,\infty)$ to measure distances.
@@ 3 No way Home: when world models and policies go out of trajectories
World models are usually learned toward minimizing a reward loss $L_{{R}}$ and/or transition loss $L_{P}$ from experiences $\eta$ collected along the agent’s trajectories. Those experiences are either gathered in the form of a batch or a replay buffer $\mathcal{B}$. In general, the loss functions take the following form: $L_{{R}}=\mathop{\mathbb{E}}_{\eta\sim\mathcal{B}}f_{{R}}\mathopen{}\mathclose{{\left(\phi,\,\overline{\!{{R}}};\,\eta}}\right)$ and $L_{P}=\mathop{\mathbb{E}}_{\eta\sim\mathcal{B}}f_{P}\mathopen{}\mathclose{{\left(\phi,\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu;\,\eta}}\right)$, where $f_{R}$ (resp. $f_{P}$) assign a “cost” relative to the error between R{R} and $\overline{\!{{R}}}$ (resp. $P$ and $\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu$) according to the experiences $\eta$ and their representation. Henceforth, we refer to the policy $\pi_{\text{b}}$ used to insert experiences in $\mathcal{B}$ as the behavioral policy.
@@ 3.1 Out-of-trajectory world model
To illustrate this problem, consider the world model of Figure 1. Assume the model is trained by collecting trajectories produced by $\pi_{\text{b}}$ in $\mathcal{M}$ where $\pi_{\text{b}}\mathopen{}\mathclose{{\left(a_{2}\mid s}}\right)\leq\epsilon$ for all $s\in\mathcal{S}_{1}$, with $\epsilon>0$. For a sufficiently small $\epsilon$, the region $\mathcal{S}_{3}$ in the original environment would remain largely unexplored while having almost no impact on the losses $L_{{R}},L_{P}$. Therefore, the representation of states in $\mathcal{S}_{3}$ ($\bar{s}_{3}$ and $\bar{s}^{\prime}_{3}$) may turn completely inaccurate. Here, the model incorrectly assigns a reward of $20$ to $\bar{s}^{\prime}_{3}$, whereas the true reward is strictly negative. Consequently, the optimal policy in $\overline{\!{\mathcal{M}}}$ deterministically selects $a_{2}$ in $\bar{s}_{1}$. When executed in the original environment, this policy drives the agent to $\mathcal{S}_{3}$ thereby degrading the behavioral policy $\pi_{\text{b}}$.
@@ 3.2 Confounding policy update
Updating both the representation and the policy solely from experience collected under a behavioral policy can degrade performance rather than improve it. In the same spirit as policy confounding (Suau et al., 2024), we call this phenomenon confounding policy update. The MDP in Figure 2 illustrates the issue.
@@
The agent maps the states $s_{2}$ and $s_{3}$ to the same latent state $\bar{s}$, i.e. $\phi(s)=\bar{s}$ iff $s\in\{s_{2},s_{3}\}$. States $s_{1}$ and $s_{4}$ each have their own latent state. We consider the behavioral policy $\pi_{\text{b}}:=\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}\circ\phi$, where $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}$ is a stochastic policy with a small exploration rate $\zeta$:
@@
for $0<\zeta\ll\epsilon$. A good representation would ideally group states from which the agent behaves similarly. Because trajectories that reach $s_{3}$ and pick $a_{2}$ are unlikely, the two states appear identical under $\pi_{\text{b}}$: $|V^{\pi_{\text{b}}}(s_{2})-V^{\pi_{\text{b}}}(s_{3})|\!\approx\!0$. Therefore, this justifies using $\phi$ as representation for $\pi_{\text{b}}$, because the values of $s_{2}$ and $s_{3}$ are nearly identical: the agent exhibits close behaviors under $\pi_{\text{b}}$ from those states.
@@
Our goal is to establish sufficient conditions to guarantee safe policy improvement during the RL process, either based on world models, state representations, or both, thus alleviating OOT world model and confounding policy update issues. Notice that, in the examples, both problems occur when performing aggressive updates from $\pi_{\text{b}}$ to a new policy $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ (the mode of the distributions drastically shifts). Intuitively, smooth updates indeed ensure to alleviate those issues: constraining the policy search to policies “close” to $\pi_{\text{b}}$ (i) prevents hallucinations in parts of the world model that have been underexplored; (ii) reduces the risk of significantly degrading the return when updating the policy. While the benefits of regularizing policy improvements have already been both theoretically and practically justified (e.g., Geist et al. 2019; Kuba et al. 2022), their implications when mixing model-based and representation learning in RL have been underexplored.
Roadmap.
To rigorously address the OOT and confounding-update issues, the next sections develop the theoretical foundations of our approach, showing how controlled policy updates, local model losses, and representation stability interact. We briefly summarize how the main results connect.
Our analysis combines neighborhood-restricted policy updates, model-quality bounds, and representation guarantees. Sect. 4 introduces the neighborhood operator defining a trust region around the behavioral policy; restricting updates to this region ensures monotonic improvement and convergence (Thm. 1). Sect. 5 then links the reward and transition losses to value discrepancies: Thm. 2 shows that, when these losses are small, and updates remain in the neighborhood, the world model stays accurate under the learned representation. Combining these ingredients yields our first SPI result (Thm. 3), guaranteeing that direct policy updates in the world model translate to improvement under controlled error. Finally, Thm. 4 shows that the same loss-based control stabilizes the encoder, ensuring that value-distinct states remain separated in the latent space.